Write the equation of a line ( y = mx + b ) given that the slope = 2/3 and the y-intercept is (0, 4).
Answer by Alan3354(69443)

Write the equation of a line ( y = mx + b ) given that the slope = 2/3 and the y-intercept is (0, 4).
------------
m = 2/3
b = 4
---------